OpenStudy (anonymous):

Write and solve a system of equations to find the cost of a large popcorn and the cost of a small drink. receipt 1 3 large popcorn 2 small drinks total due: $21.00 receipt 2 2 large popcorn 4small drinks total due: $22.00

OpenStudy (anonymous):

3p + 2d = 21 2p + 4d = 22 Solve for p and d!

OpenStudy (anonymous):

oh my gosh thank you, sorry that its such a stupid problem :(

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

3p + 2d = 21 p + 2d = 11 Subtract them!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Your "sec" takes forever, then all you said is "what" !!!

OpenStudy (anonymous):

im sorry your confusing me just nvm